1. You have observed the following returns on Stock A's stocks over the last five years:...

1.

You have observed the following returns on Stock A's stocks over the last five years:

35.30%, 27.91%, 6.08%, 38.60%, 27.03%

What is the average return on the stock over this five-year period?
Answer #1

Stock returns of stock A over the last 5 years are:

1 35.3%

2. 27.91%

3. 6.08%

4. 36.60%

5. 27.03%

thus average return on the stock over this five year period = (R1+R2+R3+R4+R5)/5

= (35.30 + 27.91 + 6.08 + 36.60 + 27.03) / 5 = 132.92 / 5 = 26.584% pa

Hence, the average return on the stock over the given five year period is 26.584% pa